# Can Large Multimodal Models Understand Agricultural Scenes? Benchmarking with AgroMind

### v1

@@ B.2 Data pre-processing details
To address the heterogeneity inherent in multi-source inputs, we propose a unified pre-processing framework characterized by both diversity and depth. The framework consists of four sequential stages—(i) multilevel annotation refinement, (ii) QA schema harmonization, (iii) logical reasoning enhancement, and (iv) task alignment—each contributing to improved data fidelity and downstream model robustness. First, given that each source employs distinct labels, formats, and pre-processing procedures, we manually refine most original annotations and a subset of images to derive secondary and tertiary annotations. For instance, pre-processing the near-infrared channel data of the Agriculture Vision Challenge[50] collection and the semantic masks from PhenoBench Dataset[51] facilitates extraction of instance-level annotations for coverage areas and individual objects. Second, QA pair generation is managed via JSON-formatted label files: each file aggregates multiple entries, with each entry linking an image, a question, and its corresponding answer, thereby ensuring high operational efficiency. Finally, through hierarchical technology integration and task-driven optimization, the pre-processing pipeline transforms raw inputs into a sophisticated, logically coherent evaluation benchmark. This approach not only addresses the challenges of format heterogeneity but also enhances data robustness via targeted augmentation strategies, providing comprehensive support for the robust training and precise evaluation of large multimodal models.
@@
2018 AI Challenge Dataset. We parsed original leaf-image annotations to extract crop species, disease category, and severity level. Using spatial disease maps, we instantiated QA templates for crop classification, health-status inference, and severity quantification. To guarantee comprehensive coverage, we applied stratified downsampling to the generated QA pairs—ensuring each crop–disease combination is represented—and then randomly sampled the balanced set to yield a diverse and complete QA corpus.
@@ C.1 Human participants
GPT-4o [11] 69.65 29.88 49.72 38.67 20.58 42.98
Gemini-1.5-Flash [12] 65.36 33.15 46.63 35.66 20.11 41.21
Claude-3.5-Sonnet [54] 46.16 15.95 40.84 30.22 18.02 33.33
DeepSeek-VL2-tiny [59] 69.79 25.49 27.49 21.71 15.74 27.51
DeepSeek-VL2-small [59] 70.16 32.46 29.03 21.39 2.47 27.84
TinyLLaVA [55] 61.30 27.05 33.70 23.40 6.75 29.88
InternVL2-2B [30] 61.23 26.80 30.69 20.50 11.56 28.49
InternVL2-4B [30] 59.60 26.32 36.88 25.15 5.34 31.46
XComposer2-4KHD [56] 63.52 25.46 36.22 30.90 11.06 32.31
LLaVA-NeXT-7B-Mistral [32] 50.59 27.61 36.38 20.60 9.43 31.07
LLaVA-NeXT-7B-Vicuna [32] 54.06 28.12 24.45 18.75 11.84 24.61
InstructBLIP-Vicuna-7B [57] 46.23 9.68 19.57 21.50 5.18 17.70
LLaVA-NeXT-Interleave-7B [31] 38.40 14.56 18.05 22.19 9.78 17.83
Mantis-LLaMA3-SigLIP [60] 56.87 24.41 36.60 24.78 11.52 31.60
Mantis-Idefics2 [60] 57.53 22.49 36.31 25.57 16.46 31.84
LLaVA-NeXT-8B [32] 69.65 28.41 36.69 21.76 6.03 32.07
InternVL2-8B [30] 71.12 29.88 39.94 29.48 6.25 34.88
Idefics-2-8b [58] 65.14 28.02 34.19 20.13 11.18 30.85
LLaVA-NeXT-13B [32] 60.56 27.85 36.98 20.39 5.12 31.43
InternVL2-26B [30] 61.52 30.15 42.48 29.85 5.81 35.84
LLaVA-NeXT-34B [32] 51.40 31.02 41.90 25.78 6.87 34.93
@@ C.2 Evaluation models and computational resource
We compared different models across multiple tasks in our benchmark to understand their capabilities in agricultural scenarios. We evaluated open-source models, including DeepSeek-VL2 series[59], TinyLLaVA[55], InternVL2 series[30], XComposer2[56], LLaVA-NeXT series[32], InstrucBLIP[57], LLaVA-NeXT-Interleave[31], Mantis series[60], and Idefics2[58], as well as closed-source models such as GPT-4o[11], Gemini[12], and Claude[54]. All experiments were conducted using an NVIDIA A800 GPU. Table 5 provides details of these models, and the specific computational resource requirements for each model are included in their respective official URLs listed in the table.
@@ C.4 More details of evaluation protocol
Open-ended questions. For open-ended questions that allow for varied expression, we use semantic similarity to evaluate correctness. Both the predicted and reference answers are embedded using a pre-trained Sentence-BERT model[61], and their cosine similarity is computed. A prediction is deemed correct if the similarity score exceeds a pre-defined threshold (e.g., 0.85), allowing for flexible yet reliable evaluation.
@@
Table 7: Performance comparison with aggregated question scenes.
(3,200) (3,200) (2,547) (3,870) (2,358) (1,800) (3,598) (2,000) (2,453) (25,026)
Human 16.04 33.36 44.76 53.43 20.63 18.98 31.46 43.82 43.47 34.72
Random 13.64 25.80 12.41 25.66 15.49 10.25 18.23 20.00 24.54 19.25
GPT-4o [11] 19.09 48.97 54.34 74.16 23.24 35.50 33.71 28.30 55.40 42.98
Gemini-1.5-Flash [12] 17.09 46.00 53.87 69.95 23.66 35.61 31.71 31.60 50.67 41.21
Claude-3.5-Sonnet [54] 14.53 43.12 29.96 49.59 16.03 34.28 33.99 28.65 41.66 33.33
DeepSeek-VL2-tiny [59] 17.31 16.66 37.77 56.54 27.14 14.06 6.78 36.80 31.59 27.51
DeepSeek-VL2-small [59] 11.06 20.44 47.51 57.39 26.12 0.22 19.84 18.70 33.43 27.84
TinyLLaVA [55] 14.00 17.47 52.30 56.82 13.61 9.56 31.88 17.45 38.73 29.88
InternVL2-2B [30] 14.09 18.50 40.40 43.85 22.52 18.94 30.85 29.35 32.25 28.49
InternVL2-4B [30] 13.34 29.56 41.30 47.49 22.90 30.78 33.07 14.65 42.07 31.46
XComposer2-4KHD [56] 13.47 24.03 38.01 47.86 24.17 44.11 32.43 25.65 41.66 32.31
LLaVA-NeXT-7B-Mistral [32] 21.56 27.22 45.78 42.95 12.60 8.50 37.99 34.35 36.00 31.07
LLaVA-NeXT-7B-Vicuna [32] 13.19 18.59 32.78 47.03 22.14 18.28 13.34 18.00 32.41 24.61
InstructBLIP-Vicuna-7B [57] 9.62 15.06 23.05 33.07 8.78 10.28 10.17 18.65 26.13 17.70
LLaVA-NeXT-Interleave-7B [31] 13.53 11.59 17.55 29.79 18.49 24.11 8.59 10.35 27.44 17.83
Mantis-LLaMA3-SigLIP [60] 21.22 24.75 45.35 57.83 21.54 33.94 16.56 21.20 36.89 31.60
Mantis-Idefics2 [60] 20.75 32.75 45.74 58.17 13.49 28.72 14.51 28.75 37.06 31.84
LLaVA-NeXT-8B [32] 17.62 23.97 46.33 49.28 19.13 25.67 37.49 16.95 41.01 32.07
InternVL2-8B [30] 16.34 29.81 40.09 51.86 27.82 33.72 35.49 25.35 47.94 34.88
Idefics-2-8b [58] 19.81 23.94 53.47 55.58 18.02 10.67 22.35 23.50 37.34 30.85
LLaVA-NeXT-13B [32] 19.31 16.84 53.28 49.92 16.37 34.61 31.16 17.25 38.52 31.43
InternVL2-26B [30] 12.72 30.53 45.78 63.49 21.97 32.11 34.44 17.35 52.22 35.84
LLaVA-NeXT-34B [32] 21.97 29.25 48.72 52.95 17.09 33.78 37.58 16.10 45.98 34.93

### v2

@@ B.2 Data pre-processing details
To address the heterogeneity inherent in multi-source inputs, we propose a unified pre-processing framework characterized by both diversity and depth. The framework consists of four sequential stages—(i) multilevel annotation refinement, (ii) QA schema harmonization, (iii) logical reasoning enhancement, and (iv) task alignment—each contributing to improved data fidelity and downstream model robustness. First, given that each source employs distinct labels, formats, and pre-processing procedures, we manually refine most original annotations and a subset of images to derive secondary and tertiary annotations. For instance, pre-processing the near-infrared channel data of the Agriculture Vision ChallengeChiu_2020_CVPR collection and the semantic masks from PhenoBench Datasetweyler2024pami facilitates extraction of instance-level annotations for coverage areas and individual objects. Second, QA pair generation is managed via JSON-formatted label files: each file aggregates multiple entries, with each entry linking an image, a question, and its corresponding answer, thereby ensuring high operational efficiency. Crucially, the logical reasoning enhancement stage embeds artificially pre-defining hierarchical reasoning processes into QA design: rather than directly using raw labels as answers, we construct step-by-step reasoning paths (e.g., for canopy diameter calculation: locate bounding box → identify tree → measure pixel coverage → convert to physical scale; for weed control: quantify coverage → analyze occlusion → generate agronomic decisions). Finally, through hierarchical technology integration and task-driven optimization, the pre-processing pipeline transforms raw inputs into a sophisticated, logically coherent evaluation benchmark. This approach not only addresses the challenges of format heterogeneity but also enhances data robustness via targeted augmentation strategies, providing comprehensive support for the robust training and precise evaluation of large multimodal models.
@@
2018 AI Challenge Dataset. We parsed original leaf-image annotations to extract crop species, disease category, and severity level, after first cleansing augmented samples by excluding flipped, rotated, and duplicated images identified through filename markers. Using spatial disease maps, we instantiated QA templates for crop classification, health-status inference, and severity quantification. To guarantee comprehensive coverage, we applied stratified downsampling to the generated QA pairs—ensuring each crop–disease combination is represented—and then randomly sampled the balanced set to yield a diverse and complete QA corpus.
@@
    Gemini-1.5-Pro     2024     N/A     https://ai.google.dev/gemini-api/docs/models/gemini
@@
GeoChat     GeoChat     2024     7B     https://huggingface.co/MBZUAI/geochat-7B
GeoLLaVA-8K     GeoLLaVA-8K     2025     7B     https://huggingface.co/initiacms/GeoLLaVA-8K
@@ C.1 Human participants
GPT-4o openai2024gpt4o 69.65 29.88 49.72 38.67 20.58 42.98
Gemini-1.5-Flash team2023gemini 65.36 33.15 46.63 35.66 20.11 41.21
Gemini-1.5-Pro team2023gemini 70.15 26.96 45.74 37.23 11.88 38.99
Claude-3.5-Sonnet anthropic2024claude3 46.16 15.95 40.84 30.22 18.02 33.33
DeepSeek-VL2-tiny deepseekvl2 69.79 25.49 27.49 21.71 15.74 27.51
DeepSeek-VL2-small deepseekvl2 70.16 32.46 29.03 21.39 2.47 27.84
TinyLLaVA TinyLLaVA 61.30 27.05 33.70 23.40 6.75 29.88
InternVL2-2B InternVL2 61.23 26.80 30.69 20.50 11.56 28.49
InternVL2-4B InternVL2 59.60 26.32 36.88 25.15 5.34 31.46
XComposer2-4KHD XComposer2 63.52 25.46 36.22 30.90 11.06 32.31
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 50.59 27.61 36.38 20.60 9.43 31.07
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 54.06 28.12 24.45 18.75 11.84 24.61
InstructBLIP-Vicuna-7B InstrucBLIP 46.23 9.68 19.57 21.50 5.18 17.70
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 38.40 14.56 18.05 22.19 9.78 17.83
Mantis-LLaMA3-SigLIP Mantis 56.87 24.41 36.60 24.78 11.52 31.60
Mantis-Idefics2 Mantis 57.53 22.49 36.31 25.57 16.46 31.84
LLaVA-NeXT-8B LLaVA-NeXT 69.65 28.41 36.69 21.76 6.03 32.07
InternVL2-8B InternVL2 71.12 29.88 39.94 29.48 6.25 34.88
Idefics-2-8b Idefics2 65.14 28.02 34.19 20.13 11.18 30.85
LLaVA-NeXT-13B LLaVA-NeXT 60.56 27.85 36.98 20.39 5.12 31.43
InternVL2-26B InternVL2 61.52 30.15 42.48 29.85 5.81 35.84
LLaVA-NeXT-34B LLaVA-NeXT 51.40 31.02 41.90 25.78 6.87 34.93
GeoChat kuckreja2023geochat 40.30 27.70 29.73 23.40 3.75 26.15
GeoLLaVA-8K wang2025geollava 53.73 0.00 18.16 21.28 1.88 15.24
@@ C.2 Evaluation models and computational resource
We compared different models across multiple tasks in our benchmark to understand their capabilities in agricultural scenarios. We evaluated open-source models, including DeepSeek-VL2 seriesdeepseekvl2 , TinyLLaVATinyLLaVA , InternVL2 seriesInternVL2 , XComposer2XComposer2 , LLaVA-NeXT seriesLLaVA-NeXT , InstrucBLIPInstrucBLIP , LLaVA-NeXT-InterleaveLLaVA-NeXT-Interleave , Mantis seriesMantis , Idefics2Idefics2 , GeoChat kuckreja2023geochat , and GeoLLaVA-8K, as well as closed-source models such as GPT-4oopenai2024gpt4o , Geminiteam2023gemini , and Claudeanthropic2024claude3 . All experiments were conducted using an NVIDIA A800 GPU. Table 5 provides details of these models, and the specific computational resource requirements for each model are included in their respective official URLs listed in the table.
Table 7: Performance comparison with aggregated question scenes.
Human 16.04 33.36 44.76 53.43 20.63 18.98 31.46 43.82 43.47 34.72
Random 13.64 25.80 12.41 25.66 15.49 10.25 18.23 20.00 24.54 19.25
GPT-4o openai2024gpt4o 19.09 48.97 54.34 74.16 23.24 35.50 33.71 28.30 55.40 42.98
Gemini-1.5-Flash team2023gemini 17.09 46.00 53.87 69.95 23.66 35.61 31.71 31.60 50.67 41.21
Gemini-1.5-Pro team2023gemini 16.25 39.38 47.64 67.53 25.32 35.00 35.65 19.50 51.24 38.99
Claude-3.5-Sonnet anthropic2024claude3 14.53 43.12 29.96 49.59 16.03 34.28 33.99 28.65 41.66 33.33
DeepSeek-VL2-tiny deepseekvl2 17.31 16.66 37.77 56.54 27.14 14.06 6.78 36.80 31.59 27.51
DeepSeek-VL2-small deepseekvl2 11.06 20.44 47.51 57.39 26.12 0.22 19.84 18.70 33.43 27.84
TinyLLaVA TinyLLaVA 14.00 17.47 52.30 56.82 13.61 9.56 31.88 17.45 38.73 29.88
InternVL2-2B InternVL2 14.09 18.50 40.40 43.85 22.52 18.94 30.85 29.35 32.25 28.49
InternVL2-4B InternVL2 13.34 29.56 41.30 47.49 22.90 30.78 33.07 14.65 42.07 31.46
XComposer2-4KHD XComposer2 13.47 24.03 38.01 47.86 24.17 44.11 32.43 25.65 41.66 32.31
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 21.56 27.22 45.78 42.95 12.60 8.50 37.99 34.35 36.00 31.07
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 13.19 18.59 32.78 47.03 22.14 18.28 13.34 18.00 32.41 24.61
InstructBLIP-Vicuna-7B InstrucBLIP 9.62 15.06 23.05 33.07 8.78 10.28 10.17 18.65 26.13 17.70
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 13.53 11.59 17.55 29.79 18.49 24.11 8.59 10.35 27.44 17.83
Mantis-LLaMA3-SigLIP Mantis 21.22 24.75 45.35 57.83 21.54 33.94 16.56 21.20 36.89 31.60
Mantis-Idefics2 Mantis 20.75 32.75 45.74 58.17 13.49 28.72 14.51 28.75 37.06 31.84
LLaVA-NeXT-8B LLaVA-NeXT 17.62 23.97 46.33 49.28 19.13 25.67 37.49 16.95 41.01 32.07
InternVL2-8B InternVL2 16.34 29.81 40.09 51.86 27.82 33.72 35.49 25.35 47.94 34.88
Idefics-2-8b Idefics2 19.81 23.94 53.47 55.58 18.02 10.67 22.35 23.50 37.34 30.85
LLaVA-NeXT-13B LLaVA-NeXT 19.31 16.84 53.28 49.92 16.37 34.61 31.16 17.25 38.52 31.43
InternVL2-26B InternVL2 12.72 30.53 45.78 63.49 21.97 32.11 34.44 17.35 52.22 35.84
LLaVA-NeXT-34B LLaVA-NeXT 21.97 29.25 48.72 52.95 17.09 33.78 37.58 16.10 45.98 34.93
GeoChat kuckreja2023geochat 12.81 16.88 41.34 43.12 10.73 36.11 21.45 17.00 35.12 26.15
GeoLLaVA-8K wang2025geollava 10.31 15.00 12.60 20.52 14.16 0.00 12.81 22.50 26.45 15.24
Table 8: Performance comparison with aggregated question sensors.
Model UAV Satellites Camera Overall
Human 42.17 28.89 34.72 34.72
Random 20.00 18.88 19.38 19.25
GPT-4o openai2024gpt4o 28.30 34.41 50.87 42.98
Gemini-1.5-Flash team2023gemini 31.60 31.68 49.00 41.21
Gemini-1.5-Pro team2023gemini 19.50 33.01 45.85 38.99
Claude-3.5-Sonnet anthropic2024claude3 28.65 29.29 36.45 33.33
DeepSeek-VL2-tiny deepseekvl2 36.80 17.00 33.22 27.51
DeepSeek-VL2-small deepseekvl2 18.70 20.41 34.16 27.84
TinyLLaVA TinyLLaVA 17.45 27.51 33.27 29.88
InternVL2-2B InternVL2 29.35 25.42 30.42 28.49
InternVL2-4B InternVL2 14.65 28.63 35.79 31.46
XComposer2-4KHD XComposer2 25.65 28.32 35.96 32.31
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 34.35 31.78 30.12 31.07
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 18.00 18.34 29.77 24.61
InstructBLIP-Vicuna-7B InstrucBLIP 18.65 14.21 19.90 17.70
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 10.35 15.30 20.62 17.83
Mantis-LLaMA3-SigLIP Mantis 21.20 23.57 38.50 31.60
Mantis-Idefics2 Mantis 28.75 22.65 38.47 31.84
LLaVA-NeXT-8B LLaVA-NeXT 16.95 31.55 34.61 32.07
InternVL2-8B InternVL2 25.35 32.17 38.08 34.88
Idefics-2-8b Idefics2 23.50 25.45 35.54 30.85
LLaVA-NeXT-13B LLaVA-NeXT 17.25 29.01 35.11 31.43
InternVL2-26B InternVL2 17.35 31.64 41.35 35.84
LLaVA-NeXT-34B LLaVA-NeXT 16.10 34.41 38.02 34.93
GeoChat kuckreja2023geochat 17.00 22.04 30.25 26.15
GeoLLaVA-8K wang2025geollava 22.50 15.53 13.99 15.24
@@ C.4 More details of evaluation protocol
Open-ended questions. For open-ended questions, we have clear and fixed reference answers. We consider a prediction correct only when it is semantically consistent with the reference answer. Therefore, we use the paraphrase-MiniLM-L6-V2 reimers2019sentence model to compute cosine similarity between predicted and reference answers. We consider an answer correct only when similarity exceeds the set threshold. Through preliminary experiments on a dataset subset, we determined the threshold as 0.60, under which all models achieve reasonable accuracy. We plan to employ expert human evaluation and calculate the overlap rate between expert and model judgments to determine the optimal threshold that maximizes agreement between expert and model.
@@ C.5 More experimental results and discussions
Table 8 illustrates the performance of different models on images captured by various sensors including unmanned aerial vehicles (UAV), satellites (SL), and cameras (CM). On UAV-captured images, where the objects are typically smaller, and the context tends to be more complex due to higher variability in altitude, angles, and resolutions, most models struggle significantly. In contrast, human participants demonstrate a notably higher understanding of these images, achieving accuracy rates up to 42.17%. However, for SL and CM images, which generally have clearer object representations and more consistent viewpoints, the performance of human evaluators is markedly lower than that of LMMs. This discrepancy indicates that LMMs excel at interpreting more structured and standardized image inputs, whereas their performance sharply declines with UAV imagery characterized by greater complexity and variability. Therefore, further enhancements in understanding UAV-captured images remain a critical area for future research and development.
@@
The observed performance imbalance across agricultural tasks has significant implications for fairness and potential biases in deploying LMMs for global agricultural applications. Given AgroMind’s extensive geographic and agricultural diversity, performance disparities in tasks like spatial localization and quantitative reasoning may disproportionately affect regions relying heavily on precision agriculture and resource optimization. Specifically, regions characterized by complex agricultural landscapes, smaller-scale farming systems, or diverse polycultures—often prevalent in parts of Africa, Southeast Asia, and Latin America—might not fully benefit from current model capabilities, exacerbating existing technological divides. Therefore, it is critical to address these imbalances through targeted model training and fine-tuning, ensuring equitable performance across different agricultural contexts. Enhancing the robustness and adaptability of LMMs in challenging scenarios, such as UAV-captured imagery, will be essential to prevent bias and promote fair access to advanced agricultural technologies worldwide.